x 0 1 2 3 4 5
P(Xx) 0.18 0.35 0.47 0.73 0.88 1

a. Calculate P(X ≤ 1). (Round your answer to 2 decimal places.)




b. Calculate P(X = 3). (Round your answer to 2 decimal places.)



c. Calculate P(1 ≤ X ≤ 3). (Round your answer to 2 decimal places.)
Answer #1

a)

P(X < 1) 0.35

We get directly from given tables.

b)

P(X=3)=P(X\leq 3)-P(X\leq 2)

We can get above cumulative probabilities from given table. So,

P(X = 3) = 0.73-0.47 = 0.26

c)

P(1\leq X\leq 3)=P(X\leq 3)-P(X\leq 0)

We can get above cumulative probabilities from given table. So,

P(I < X < 3) 0.73-0.18-0.55